OpenStudy (anonymous):

How do I get the derivative of ln(x^2)?

OpenStudy (anonymous):

2/x

OpenStudy (anonymous):

why though? isn't it the chain rule?

OpenStudy (anonymous):

yes. d(ln u)/dx = 1/u(du/dx). So [1/(x^2)]*2x which simplifies to 2/x.

OpenStudy (anonymous):

outer function: ln() inner function: x^2 ... d(outer function) = 1/x d(inner function) = 2x 1/x * x^2 * 2x = 2x^2 ???

OpenStudy (anonymous):

the d(outer function)= 1/(x^2) not 1/x.